# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v3

@@ G.3 Deep architectures
In App. E.1, we discussed how to derive $p_{f_{0}}\left({x}\right)$ for a neural network with a single hidden layer. In the following section, we will discuss the extension of the computation to a network with an arbitrary number of hidden layers. After that, we will focus on the impact that depth has on IGB. In particular, in App. G.3.2 we will show how, in the presence of IGB, this can be amplified as the network’s depth increases.
@@ G.3.1 Multi-layer perceptron
The strategy that we will follow is similar to the one presented for the single hidden layer counterpart (see App. E.1). In other word we will propagate across the network layers keeping track of the changes in the distributions $p_{h^{(l+1)}_{i}}^{(\chi)}\left({x}\right)$ and $p_{\left\langle{h^{(l+1)}_{i}}\right\rangle}\left({x}\right)$. Using CLT and its extentions considerably simplifies this propagation process because to keep track of changes in the distribution it is sufficient to see how certain quantities vary 131313a Gaussian distribution is completely determined from the first two cumulants. Compared to the analysis in App. E.1 however, there is an important complicating element. While the elements in the set $\{h^{(1)}_{i}\}$ follow the same distribution (regardless of the choice of activation function), from the second layer onward the activation function can induce a break in symmetry among the layer nodes, in the sense that they will follow, fixed a given configuration for the network weights, different distributions. This symmetry breaking, as we shall see, can cause an accentuation of the IGB effect. To show this point we will consider in our analysis the ReLU activation function 141414we do not include the MaxPooling to underlining the amplification of the IGB; in the presence of MaxPooling, increasing the kernel size we observe strong IGB even with a single hidden layer architectures. Note that absence of MaxPooling is equivalent of a MaxPooling with minimum kernel size. Starting from the same setting described in App. E.1 we will derive a set of iterative equation to propagate across multiple layers.
@@
    as showed in App. E.1 and in App. D.2, in the limit $d\rightarrow\infty$,
@@
    All the nodes $\{h^{(1)}_{i}\}$ follow the same distribution and so do the nodes $\{g^{(1)}_{i}\}$ . In particular, this means that
@@
    It is easy to proof that the generic r.v. $\left(w^{(1)}_{ij}g^{(1)}_{j}\right)$ involved in the sum satisfies the conditions of Eq. (29) and Eq. (30). So, in the limit $N_{1}\rightarrow\infty$, again we have a convergence to a normal distribution. In particular
@@
    In the last step of Eq. (166), as done for $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ in Sec. E.1, we used the concentration result derived in App. C.2.1. In particular that the distribution of the r.v. $S_{\boldsymbol{w}_{i}}\equiv\sum_{j=1}^{N_{1}}w^{(1)}_{ij}$ stays stable in the limit $N_{1}\rightarrow\infty$ 151515fluctuations and mean stay both $\mathcal{O}\left(1\right)$, while the distribution of $S_{\boldsymbol{w}_{i}^{2}}\equiv\sum_{j=1}^{N_{1}}\left(w^{(1)}_{ij}\right)^{2}$ asymptotically narrow around the mean value. Note that, as the distribution of $S_{\boldsymbol{w}_{i}}$ does not asymptotically concentrate, and since $\left\langle{g^{(1)}}\right\rangle\neq 0$, each node in the set $\{h^{(2)}_{i}\}$ will follows a different distribution. In particular we will have a set of normally distributed r.v.s, centred in different random points, $\{\left\langle{g^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}}\}_{i}$. After passing through the ReLU activation function we will have
@@
    We can repeat the approach of the previous layer for the new set of variables
@@
  • layer-l
    The steps described for the propagation of the layer-3 provide us an iteration scheme that we can follow for a generic layer $l\geq 3$. In fact, knowing the statistics of the previous layers, we can compute
@@
Figure G.1: Effects induced by the depth of the network. Simulations on an MLP with $L$ hidden layers and ReLU activation function (without any pooling layers) are compared for two different depth values. The theoretical curves are superimposed on the empirical distributions, demonstrating a good agreement between the predictions of our analysis and empirical observations. For a single hidden layer network $(L=1)$ we have a stable distribution (i.e. that does not narrow around $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$ for $D\rightarrow\infty$) but still peaked on $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$. Increasing the number of hidden layers the probability mass moves from the center to the extremes of the support. For these simulations we used the GB dataset on model MHLP (see App. I.3 for more details).
G.3.2 Amplification of IGB with depth
Here we prove that the distribution of $\mathbb{P}^{\mathcal{W}}\left(f_{0}\right)$ converge to a delta peaked distribution in a multi-layer perceptron with ReLU activation when the number of layers goes to infinity. More precisely
@@ G.3.2 Amplification of IGB with depth
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP, mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. Let us consider a multi-layer perceptron with $L$ hidden layers, ReLU activation function and without pooling layers. Let us assume that, in the $\lim_{N_{l}\rightarrow\infty}\forall l\in[0,\dots,L]$, w.h.p.: $\textrm{Var}_{\chi}\left({h^{(l)}}\right)>0,\;\forall l\,$, (180) and in particular $\lim_{l\rightarrow\infty}\textrm{Var}_{\chi}\left({h^{(l)}}\right)>0,$ (181) where $\textrm{Var}_{\chi}\left({h^{(l)}}\right)$ indicate the variance of $\mathbb{P}^{\chi}\left(h^{(l)}\mid\mathcal{W}\right)$. Then $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty\,.$ (182)
To prove (182) we will show that
@@
$\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(l+2)}}\right)}>\left(\frac{1}{K}\right)^{l}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\left(K+\epsilon\right)^{l}>l\underbrace{\left(\frac{1}{K}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\epsilon\right)}_{\equiv\frac{1}{C}}$
@@
Let us start from Eq. (183). From Eq. (175)
@@
Note that $\mu_{+}\mu_{-}=0$ implies either $\mu_{+}=0$ or $\mu_{-}=0$ i.e. a distribution $p_{h^{(l)}_{j}}^{(\chi)}\left({x}\right)$ with non-positive or non-negative support. Since $p_{h^{(l)}_{j}}^{(\chi)}\left({x}\right)$ is a Gaussian distribution this may only happen in the limit of its variance going to 0 161616i.e., if the Gaussian shrink into a Dirac delta distribution. By hypothesis we excluded this possibility (Eq. (180) and Eq. (181)); therefore $\mu_{+}\mu_{-}<0$ and we can rewrite
@@
where for the second step we used again Eq. (189). To summarize we showed that
@@
Now let us pass to another quantity:
@@
We observe a fundamental distinction compared to the previous cases. Network depth can amplify the level of IGB in systems where it is already present. However, depth alone does not induce IGB. Considering a model with a linear activation function and no max-pooling after the first hidden layer, and using Gaussian data centered at zero, we have $\left\langle{\rho^{(1;m)}}\right\rangle=\left\langle{h^{(1)}_{m}{}}\right\rangle=0$. Thus, there is no Node Symmetry Breaking (NSB) in the second layer; in other words, the nodes in the second layer are Gaussianr.v.s centered at zero. It can be easily shown that, by iterating this process through each layer, the same scenario will be repeated in each subsequent layer up to the output layer. Therefore, NSB does not occur even in the output nodes, resulting in the absence of IGB in the system.
@@ Appendix H Extension to multi-class problems
We want now to extend the analysis to the multi-class case, i.e. to problems with $N_{C}>2$. Following the framework we introduced earlier, we can easily extend the computation. We will again consider the distribution among the $N_{C}$ classes of the dataset elements after initialization. In particular, we can define $N_{C}$ values associated with the fraction of dataset elements classified as belonging to each of the $N_{C}$ classes, $\left\{f_{i}\right\}_{i=0}^{N_{C}-1}$. We could also consider the distribution of the sorted frequencies; in other word, in each experiment, we order the classes according to the corresponding frequencies and not by their label. To distinguish these frequencies (and their statistics) from the set $\left\{f_{i}\right\}_{i=0}^{N_{C}-1}$ , we call this new set $\left\{f_{\tilde{i}}\right\}_{\tilde{i}=0}^{N_{C}-1}$. In particular $f_{\tilde{0}}$ indicates the bigger frequency among the $N_{C}$, $f_{\tilde{1}}$ the second one and so on. Finally we add the information of the set (or subset) cardinality, $M$, on the output nodes set $\left\{O^{(c)}_{M}\right\}_{c=0}^{M-1}$.  [^17] The statistics of the r.v.change with the number of classes (or nodes) Analogously to the set of fractions, we define the set of ranked output nodes $\left\{O^{(\tilde{c})}_{M}\right\}_{\tilde{c}=0}^{M-1}$ In the following, similarly to Sec. E.1 we will derive the distribution of $f_{0}$.
Figure H.1: $p_{f_{0}}\left({x}\right)$ in a multi-class problem with $N_{C}=10$. The empirical mean value is reported; note that despite the difference in the distribution the mean value is the same that we would observe using a linear activation function $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{N_{C}}$. For these simulations we used the dataset GB on the model SHLP with MaxPooling ($m=20$). (see App. I.3 for more details).

### v4

@@ G.3 Deep architectures
In App. E.1, we discussed how to derive $p_{f_{0}}\left({x}\right)$ for a neural network with a single hidden layer. In this section, we will discuss the extension of the computation to a network with an arbitrary number of hidden layers. After that, we will focus on the impact that depth has on IGB. In particular, in App. G.3.3 we will show that network depth does not cause IGB, but it amplifies it when the network’s depth increases.
@@ G.3.1 Multi-layer perceptron
The strategy that we will follow is similar to the one presented for the single-hidden-layer counterpart (see App. E.1). We will discuss here the case of MLP with ReLU activation function to focus on the effects of network depth. Extending the results discussed in the single-layer case to deeper networks, including variations such as the inclusion of Max-Pooling, is straightforward. We will propagate the signal across the network layers, keeping track of the changes in the distributions $p_{h^{(l+1)}_{i}}^{(\chi)}\left({x}\right)$ and $p_{\left\langle{h^{(l+1)}_{i}}\right\rangle}\left({x}\right)$. Using the CLT and its extensions considerably simplifies this propagation process because to keep track of changes in the distribution it is sufficient to see how few quantities vary. [^15] This is because a Gaussian distribution is completely determined by the first two cumulants. Compared to the analysis in App. E.1, however, there is an important complicating element. While the elements in the set $\{h^{(1)}_{i}\}$ follow the same distribution (regardless of the choice of the activation function), from the second layer onward the activation function can induce a breaking in symmetry among the layer nodes, in the sense that they will follow, fixed a given configuration for the network weights, different distributions. This symmetry breaking, as we shall see, can cause an accentuation of IGB. To show this point, we will consider in our analysis the ReLU activation function. [^16] We do not include the max pooling to isolate the effect of depth from that of max pooling. Note that absence of max pooling is equivalent of a max pooling with minimum kernel size. Starting from the same setting described in App. E.1, we will derive a set of iterative equations to propagate across multiple layers.
@@
    As shown in App. E.1 and in App. D.2, in the limit $d\rightarrow\infty$,
@@
    All the nodes $\{h^{(1)}_{i}\}$ follow the same distribution, and so do the nodes $\{g^{(1)}_{i}\}$ . In particular, this means that
@@
    It is easy to prove that the generic r.v. $\left(w^{(1)}_{ij}g^{(1)}_{j}\right)$ involved in the sum satisfies the conditions of Eq. (29) and Eq. (30). So, in the limit $N_{1}\rightarrow\infty$, we again have a convergence to a normal distribution. In particular,
@@
    In the last step of Eq. (167), as done for $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ in Sec. E.1, we used the concentration result derived in App. C.2.1. In particular that the distribution of the r.v. $S_{\boldsymbol{w}_{i}}\equiv\sum_{j=1}^{N_{1}}w^{(1)}_{ij}$ stays stable in the limit $N_{1}\rightarrow\infty$, [^17] Fluctuations and mean stay both $\mathcal{O}\left(1\right)$. while the distribution of $S_{\boldsymbol{w}_{i}^{2}}\equiv\sum_{j=1}^{N_{1}}\left(w^{(1)}_{ij}\right)^{2}$ asymptotically narrows around the mean value. Note that, as the distribution of $S_{\boldsymbol{w}_{i}}$ does not asymptotically concentrate, and since $\left\langle{g^{(1)}}\right\rangle\neq 0$, each node in the set $\{h^{(2)}_{i}\}$ will follow a different distribution. In particular we will have a set of normally distributed r.v.s, centred on different random points, $\{\left\langle{g^{(1)}}\right\rangle S_{\boldsymbol{w}_{i}}\}_{i}$. After passing through the ReLU activation function, we will have
@@
    We can repeat the approach of the previous layer for the new set of variables,
@@
  • layer-$\boldsymbol{l}$
    The steps described for the propagation of the layer-3 provide us with an iteration scheme that we can follow for a generic layer $l\geq 3$. In fact, knowing the statistics of the previous layers, we can compute
@@
Figure G.1: Effects induced by the depth of the network. Simulations on an MLP with $L$ hidden layers and ReLU activation function (without any pooling layers) are compared for two different depth values. The theoretical curves are superimposed on the empirical distributions, demonstrating a good agreement between the predictions of our analysis and empirical observations. For a single hidden layer network $(L=1)$ we have a stable distribution (i.e. that does not narrow around $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$ for $D\rightarrow\infty$) but still peaked on $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=0.5$. Increasing the number of hidden layers the probability mass moves from the center to the extremes of the support. For these simulations we used the GB dataset on model MHLP (see App. I.4 for more details).
G.3.2 Inclusion of Non-Null Biases
In the analysis presented in App. G.3.1, we considered a setting with null biases for the sake of clarity. However, the analysis can be extended to initializations where the biases are set to non-null values. For instance, considering an MLP with Gaussian-initialized bias parameters, the pre-activations of a generic hidden layer can be described as:
$h^{\prime}_{i}=h_{i}+b_{i}\,.$ (181)
Here, $h_{i}$ represents node $i$’s pre-activation in the absence of biases, which, as per our analysis, follows a normal distribution. [^18] We omitted the layer index $l$ on the variables $h_{i},\,b_{i},\,h^{\prime}_{i}$ to lighten the notation. The term $b_{i}$ denotes the bias parameter. Given that $h^{\prime}_{i}$ is a sum of two independent Gaussian random variables, it too will exhibit a normal distribution. This allows for a seamless extension of our analysis. Similarly, if the bias parameters are initialized with a constant value ($b_{i}=c$), the analysis can be adapted accordingly. Adding a constant to $h_{i}$ merely shifts the distribution’s center, preserving the Gaussian profile for $h^{\prime}_{i}$ and reaffirming the robustness of our findings regarding IGB across different initialization strategies.
G.3.3 Amplification of IGB with depth
Here we prove that the distribution of $\mathbb{P}^{\mathcal{W}}\left(f_{0}\right)$ converges to a delta-peaked distribution in a multi-layer perceptron with ReLU activation when the number of layers goes to infinity. More precisely:
@@ G.3.3 Amplification of IGB with depth
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP, mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. Let us consider a multi-layer perceptron with $L$ hidden layers, ReLU activation function, and without pooling layers. Let us assume that, in the $\lim_{N_{l}\rightarrow\infty}\forall l\in[0,\dots,L]$, w.h.p.: $\textrm{Var}_{\chi}\left({h^{(l)}}\right)>0,\;\forall l\,$, (182) and in particular $\lim_{l\rightarrow\infty}\textrm{Var}_{\chi}\left({h^{(l)}}\right)>0,$ (183) where $\textrm{Var}_{\chi}\left({h^{(l)}}\right)$ indicate the variance of $\mathbb{P}^{\chi}\left(h^{(l)}\mid\mathcal{W}\right)$. Then $\lim_{L\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty\,.$ (184)
To prove Eq. (184), we will show that
@@
$\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(l+2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(l+2)}}\right)}>\left(\frac{1}{K}\right)^{l}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\left(K+\epsilon\right)^{l}>l\underbrace{\left(\frac{1}{K}\frac{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(2)}}\right\rangle}\right)}{\textrm{Var}_{\chi}\left({h^{(2)}}\right)}\epsilon\right)}_{\equiv\frac{1}{C}}\,$,
@@
Let us start from Eq. (185). As a first step we prove that, for $N_{l-1}\penalty 10000\ \rightarrow\penalty 10000\ \infty$,
$\frac{1}{2}\textrm{Var}_{\chi}\left({h^{(l)}}\right)>\mkern 1.5mu\overline{\mkern-1.5mu\textrm{Var}_{\chi}\left({g^{(l)}}\right)\mkern-1.5mu}\mkern 1.5mu$ $\equiv\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu.$ (189)
To prove this inequality we will now focus separately on the two terms of the r.h.s.. First we analyze $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{g^{(l)}}\right\rangle^{2}\mkern-1.5mu}\mkern 1.5mu$. From Eq. (176)
@@
Note that $\mu_{+}\mu_{-}=0$ implies either $\mu_{+}=0$ or $\mu_{-}=0$ i.e. a distribution $p_{h^{(l)}_{j}}^{(\chi)}\left({x}\right)$ with non-positive or non-negative support. Since $p_{h^{(l)}_{j}}^{(\chi)}\left({x}\right)$ is a Gaussian distribution this may only happen in the limit of its variance going to 0 i.e. if the Gaussian shrink into a Dirac delta distribution. By hypothesis we excluded this possibility (Eq. (182) and Eq. (183)); therefore $\mu_{+}\mu_{-}<0$ and we can rewrite
@@
where for the second step we used again Eq. (192). In summary, we showed that
@@
We now turn our attention to the second term appearing in the r.h.s. of Eq. (189), i.e. $\mkern 1.5mu\overline{\mkern-1.5mu\left\langle{\left(g^{(l)}\right)^{2}}\right\rangle\mkern-1.5mu}\mkern 1.5mu$:
@@
We observe a fundamental distinction compared to the previous cases. Network depth can amplify the level of IGB in systems where it is already present. However, depth alone does not induce IGB. Considering a model with a linear activation function and no max pooling after the first hidden layer, and using Gaussian data centered at zero, we have $\left\langle{\rho^{(1;m)}}\right\rangle=\left\langle{h^{(1)}_{m}{}}\right\rangle=0$. Thus, there is no Node Symmetry Breaking (NSB) in the second layer; in other words, the nodes in the second layer are Gaussian r.v.s centered at zero. It can be easily shown that, by iterating this process through each layer, the same scenario will be repeated in each subsequent layer up to the output layer. Therefore, NSB does not occur even in the output nodes, resulting in the absence of IGB in the system.
@@ Appendix H Extension to multi-class problems
We want now to extend the analysis to the multi-class case, i.e. to problems with $N_{C}>2$. Following the framework we introduced earlier, we can easily extend the computation. We will again consider the distribution among the $N_{C}$ classes of the dataset elements after initialization. In particular, we can define $N_{C}$ values associated with the fraction of dataset elements classified as belonging to each of the $N_{C}$ classes, $\left\{f_{i}\right\}_{i=0}^{N_{C}-1}$. We could also consider the distribution of the sorted frequencies; in other words, in each experiment, we order the classes according to the corresponding frequencies and not by their label. To distinguish these frequencies (and their statistics) from the set $\left\{f_{i}\right\}_{i=0}^{N_{C}-1}$ , we use indices $\tilde{i}$, and call this new set $\left\{f_{\tilde{i}}\right\}_{\tilde{i}=0}^{N_{C}-1}$. In particular $f_{\tilde{0}}$ indicates the biggest frequency among the $N_{C}$, $f_{\tilde{1}}$ the second one and so on. Finally we add the set (or subset) cardinality, $M$, on the output nodes set $\left\{O^{(c)}_{M}\right\}_{c=0}^{M-1}$. [^19] Here, the cardinality of the set indicates the number of classes over which the statistics are computed. The statistics of the r.v. change with the number of classes (or nodes), i.e., with the cardinality. Analogously to the set of fractions, we define the set of ranked output nodes $\left\{O^{(\tilde{c})}_{M}\right\}_{\tilde{c}=0}^{M-1}$ In the following, similarly to Sec. E.1 we will derive the distribution of $f_{0}$.
Figure H.1: The distribution $p_{f_{0}}\left({x}\right)$ in a multi-class problem with $N_{C}=10$. The empirical mean value is reported; note that despite the difference in the distribution the mean value is the same that we would observe using a linear activation function $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{N_{C}}$. For these simulations we used the dataset GB on the model SHLP with max pooling ($m=20$). (see App. I.4 for more details).
@@
H.1 Increasing number of classes exacerbates IGB
Having a large number of classes increases the probability that the center of the distribution related to one of these classes is an outlier of the distribution $p_{\mu_{c}}(x)$. However, since $p_{\mu_{c}}(x)$ is Gaussian (a fast-decaying distribution), the typical value of the maximum increases slowly. Therefore, an extremely large number of classes, $N_{c}$, will be needed to observe significant changes.
We can analyze this quantitatively. The typical value of the maximum out of $N_{c}$ Gaussian random variables of variance $\epsilon$ grows as $\sqrt{(2\epsilon\log(N_{c}))}$ [Gumbel, 1958, Hartarsky et al., 2019]. In the absence of IGB, $\epsilon$ is the variance of $p_{\mu_{c}}(x)$ and scales inversely with the dataset size, i.e., $\sim 1/D$ (see e.g., App. D.2 and Fig. D.1). This implies that outliers will only appear if the number of classes far exceeds the number of datapoints, a situation that does not occur in real life.
In summary, if there is no IGB, a large enough dataset ensures that increasing the number of classes does not affect the presence of IGB.
Conversely, if IGB is present, a sufficiently large number of classes may result in high imbalance even with a relatively small value of $\gamma(\mathcal{A},\psi(\chi))$.